Station Messages
24 Sep 2018 18:04 CDT 2018267+18:04 CDT

Overview Datums Bench Marks Data Query Site Visit Daily QC Messages Attributes

138: Packery Current Meter #2
Packery Channel, Texas
27° 37' 22" N   97° 12' 55" W
PCM2
USACE - The Coastal & Hydraulics Laboratory
USACE - Coastal Inlets Research Program
United States Army Corps of Engineers

2009282+1841 (note, ndurham)
1990001+0000 (01/01/1990 00:00 - 01/01/1990 00:01) :

Positive X values are flood, negative X values are ebb.

2016137+1151 (retired, lloyd)
ALL : 2016123+0100 (05/02/2016 01:00 - 05/02/2016 01:01) :

Little data dissemination starting 2016123+0100

2016102+1309 (retired, lloyd)
2016098+0754 (04/07/2016 07:54 - 04/07/2016 07:55) :

Little data dissemination starting 2016098+0754

2016048+1421 (retired, blorentson)

No data

2016046+1431 (retired, blorentson)

No data

2016041+1454 (retired, jrizzo)
ALL : 2016039+0424 (02/08/2016 04:24 - 02/08/2016 04:25) :

Little data recovery starting 2016039+0424

2016041+1453 (retired, zhasdorff)

low data recovery

2016030+1406 (retired, jrizzo)
ALL : 2016028+0306 (01/28/2016 03:06 - 01/28/2016 03:07) :

Data dissemination sporadic starting 2016028+0306

2016022+2115 (retired, jrizzo)

sporadic current data

2016009+1945 (retired, lloyd)

No data

2015349+1452 (retired, hmahlke)

no data

2015269+1010 (retired, jrizzo)

No data

2015207+1311 (retired, jrizzo)

continuous data gaps

2015187+1336 (retired, lloyd)

No data

2015004+1417 (retired, jrizzo)

Battery @11.6V

2014346+1201 (retired, amendoza)

No data

2014338+1148 (retired, amendoza)

No data

2014322+1147 (retired, amendoza)

No data being disseminated. IP modem suspect

2014301+1150 (retired, amendoza)

No battery data

2014296+1313 (retired, amendoza)

No data

2014280+1404 (retired, lloyd)

No data

2014153+1424 (retired, lloyd)

No data

2012342+1043 (retired, ussery)

Problem with sontek; drastic difference in signal strength. Faulty current meter pulled.

2012317+2116 (retired, lloyd)

No data since 11-12-2012.

2010043+1851 (retired, ussery)

Battery voltage at 10 V.

2010008+1138 (retired, ussery)

Little data received since 2010004+1400. Field crew investigating.

2009356+0341 (retired, ussery)

Battery @7.8V, no data

2009341+1241 (retired, ussery)

No data since 2009338. Data has resumed.

2009323+1119 (retired, ussery)

No data received since 2009322+0700.

2008287+2040 (retired, ussery)

Removed current meter due to faulty cable and rotting mount. Will redeploy 10-13-2008. Sontek redeployed.

2008225+1715 (retired, ussery)

No data since 2008205+1130. Instrument reset.

2008121+1244 (retired, ussery)

No data since 2008118+1930. Instrument replaced.

2008002+1850 (retired, ussery)

Sensor malfunction. Sensor mount broken. Sensor mount repaired as per field crew.

2007269+2114 (retired, ussery)

No data from station since 2007258+0930. Sensor has malfunctioned.

2007202+2002 (retired, jdavis)

Battery voltage not reported since 2007185. -- fixed

2007089+1704 (retired, jdavis)

Data sporadic since 2007088+2230. -- The ip modem poll of gbra1 was interfering with PCM2 (writing to the same file at the same time).

2007026+1026 (retired, jrizzo)

No data since 2007025+0300. Battery replaced, data resumed

2006335+1521 (retired, jrizzo)

X & Y split. The Y component returned to normal.

2006277+1540 (retired, duff)

No data since 2006270. ADCP was in a weird mode where it wouldn't respond to the normal commands. Finally got a prompt and restarted the system. Appears to be working now.

2006035+0710 (retired, duff)
all : 12.22.2005 (12/22/2005 00:00 - 12/23/2005 00:00) :

claned and painter sensor, radio link spotty

2006034+2103 (retired, dwilliams)
vlx :

Positive X values are ebb current, negative values are flood current.

2006034+2102 (retired, dwilliams)
2006034+2059 (retired, dwilliams)
02/03/2006 (02/03/2006 00:00 - 02/04/2006 00:00) :

Instrument firmware updated allowing positive values to represent flood tide and negative values to represent ebb tide. Sign inversion applied to previous data.

2006004+1406 (retired, kjenkin)

Data recovery very spotty. Repositioned antenna on NRC roof.

2005349+1224 (retired, jrizzo)

No data since 2005348. Data resumed but spotty

2005268+1332 (retired, kjenkin)

No data since 2005265.